DuPONT AND ROE

A firm has a profit margin of 7% and an equity multiplier of 2.1. Its sales are $300 million, and it has total assets of $180 million. What is its ROE? Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1

Total asset turnover=Sales/Total asset

=300/180=1.67

ROE=Profit margin*Total asset turnover*equity multiplier

=7*1.67*2.1

=24.5%